# Accretion tori around rotating neutron stars

### v1

@@ 6 Frequency behaviour
Figure 1: The axisymmetric oscillation frequencies calculated for different $\beta$ and NSs with $j=0.2$, and two particular values of $q/j^{2}$, $q/j^{2}=1.5$ (extreme NS compactness), and $q/j^{2}=10$ (large NS oblateness). Left: The radial mode. Right: The vertical mode. The curves indicating particular values of $\beta$ are normalised to the maximal value $\beta=\beta_{\mathrm{max}}$. The critical radius limiting existence of cusp tori is denoted by the dotted vertical line marked as $r_{\mathrm{mc}}$. The maximal value $\beta=\beta_{\mathrm{max}}$ matches the cusp value when $r<r_{\mathrm{mc}}$, and the value implying tori extending to infinity when $r>r_{\mathrm{mc}}$. For the absolute value of $\beta$, only one specific curve is drawn in each panel, marked in grey, $\beta=\sqrt{0.15}$, which makes it possible to distinguish frequencies that correspond to relatively large tori, as they are displayed below this curve. The dotted vertical lines in each panel marked as $r_{mc}$ denote the radii limiting the existence of cusp tori – any torus with a centre above this radii does not exhibit the cusp. An analogous notation is later used in several other figures.
@@ 6.1 Axisymmetric modes
The left panel of Figure 1 provides a comparison between the axisymmetric ($m=0$) radial mode frequencies for tori with different thickness and NSs with $j=0.2$ and different compactness. The frequencies behave qualitatively the same as the test particle’s radial epicyclic frequency, being zero at ISCO, rising rather quickly with increasing radius to a distinct maximum and then slowly almost steadily decreasing. Because the pressure correction $\nu_{\mathrm{R}}^{(2)}$ is negative, the pressure forces tend to slow the epicyclic motion, and $\beta_{\mathrm{cusp}}$ rises with $r_{0}$, the maximum is located closer to the star than it would be in the case of a test particle epicyclic motion. A more detailed quantitative comparison is provided in Appendix A, Figure 7.
In a full analogy to the left panel of Figure 1, the right panel provides comparison between the axisymmetric ($m=0$) vertical mode frequencies for tori with different thickness and NSs with $j=0.2$ and different compactness. Again, the frequencies behave qualitatively like a test particle’s vertical epicyclic frequency while the pressure corrections are negative. A more detailed quantitative comparison is also provided in Appendix A, Figure 8.
@@ 6.2 Radial precession ($m=-1$ radial mode)
The left panel of Figure 2 provides a comparison between the radial precession mode frequencies for tori with different thicknesses and NSs with $j=0.2$ and different compactness. For the radii close to ISCO, the pressure corrections are negative. The maximal precession frequency is achieved at $r_{0}=r_{\mathrm{ISCO}}$, where only infinitely slender torus with zero pressure corrections can exist. The maximum achievable frequency slightly drops with rising NS quadrupole moment as the ISCO moves to higher radii, but it is still higher than the precession rate for lower $q$ at the same (coordinate) radius.
For more distant radii, the corrections are smaller but can be positive, and at radii where they change the sign, the frequency does not depend on the torus thickness. This feature is present for both rotating and non-rotating stars. Radii relevant to this effect are shown in the right panel of Figure 2. A more detailed quantitative comparison between the radial precession mode frequencies is provided in Appendix B, Figure 9.
@@ 6.3 Vertical precession ($m=-1$ vertical mode)
The left panel of Figure 3 provides a comparison between the vertical precession mode frequencies for tori with different thicknesses and NSs with $j=0.2$ and different compactness. The frequency behaviour is even more complex than in the case of radial precession.
For compact stars, the maximal frequencies are achieved at the ISCO. For the test-particle motion and large NS oblateness, the combined influence of the oblate shape of the star and frame-dragging induce non-monotonic frequency profiles with maximum above the ISCO and emergence of a zero-frequency radius where the precession changes its orientation (Kluźniak & Rosińska, 2013, 2014; Tsang & Pappas, 2016; Urbancová et al., 2019). The same happens for non-slender tori. At the same time, the correction to the mode frequency exhibits a behaviour similar to the case of radial precession, and radii, where the correction changes its sign exists. The relevant radii are shown in the right panel of Figure 3.
A more detailed quantitative comparison between the vertical precession mode frequencies is provided in Appendix B (Figure 10), while an extended discussion on the mode frequency behaviour is given in Section 8.3.
@@ 7 Approximate relations
The formulae for the oscillation and precession frequencies investigated in the previous section are presented in our public codes. Their explicit forms are rather lengthy and potentially too complicated for practical calculations. Inspired by our previous work (Török et al., 2022), we provide their simplified implicit versions that enable more efficient implementation.
For each of the four modes, we present these formulae corresponding to the geodesic limit of a slender torus (which coincides with those describing free test particle motion) and the limit of a marginally overflowing (cusp) torus. More general formulae are also included to enable performing calculations of arbitrarily thick tori.
@@ 7.1 Axisymmetric radial mode
we obtain a relation for the axisymmetric radial oscillations of a torus with a cusp, and $\mathcal{R}$ takes values between 0 and 1. Its dependence on the spacetime parameters is illustrated in Figure 4a while a comparison between the (exact) mode frequencies and the approximate relation is shown in Figure 4b assuming very slowly rotating stars. A thorough comparison between the outcomes of exact calculations and the application of the approximate relation across the whole considered range of parameters is then provided in Appendix C (Figure 11).
@@ 7.3 Radial precession (non-axisymmetric radial mode)
The radial precession frequency can be approximated as
@@
Choosing a constant $\mathcal{B}$, $\mathcal{B}=1$, our relation with a very high accuracy provides the value of the (geodesic) periastron precession frequency of oblate neutron stars.
@@
the resulting relation well matches the numerical findings for the cusp torus’s radial precession. The specific $\mathcal{B}$ function does not depend on the Keplerian frequency (radial coordinate $r$), unlike the case of the axisymmetric modes. It takes values close to $\mathcal{B}=0.75$. The exact values are illustrated in Figure 4e while a comparison between the (exact) mode frequencies and the approximate relation is shown in Figure 4f assuming very slowly rotating stars. A thorough comparison between the outcomes of the exact calculations and the application of the approximate relation across the whole considered range of parameters is then provided in Appendix C (Figure 13).
@@ 7.4 Vertical precession
$-\frac{9}{\nu_{0}}(q-j^{2})(\mathcal{V}_{0}^{-2/3}-7.2),$
we can describe the vertical precession of a cusp torus. The $\mathcal{T}$ function takes values very close to 1. The exact values are illustrated in Figure 4g while a comparison between the (exact) mode frequencies and the approximate relation is shown in Figure 4h assuming very slowly rotating stars. A thorough comparison between the outcomes of the exact calculations and the application of the approximate relation across the whole considered range of parameters is then provided in Appendix C (Figure 14).
@@ 7.4.2 Tori of arbitrary thickness
To describe a torus of arbitrary thickness we use
@@ 8 Discussion and conclusions
The impact of the torus thickness in most cases leads to relatively high negative corrections to the modes’ frequencies, which should be reflected within possible astrophysical applications. Furthermore, as found in Sections 6.2 and 6.3, for the precession modes, the pressure corrections can be positive within a limited range of radii because the special radii exist where the precession frequency does not depend on the torus thickness. This can have observable consequences. Nevertheless, since we are constrained by the adopted perturbative approach and the effect manifests itself on a relatively large radius where the scatter between the frequencies calculated for different tori is generally very small, it will be useful to verify its presence by a suitable future numerical solution of Equation (8).
@@ 8.1 Influence of NS parameters
The NS rotation and oblateness have a significant impact on the frequencies of the investigated tori oscillations calculated for a given torus thickness and one may clearly see the impact of the quadratic relations behind the nature of the modes. This impact can be well illustrated in terms of quadrupole parameter $q/j^{2}$, which for an NS with a given mass does not depend on the NS rotational frequency. The consequences of frame dragging caused by the NS’s angular momentum predominate at low values of $q/j^{2}$. On the other hand, at large values, the effects of frame dragging are often compensated by the (rotationally induced) NS’s oblateness. As a result, for low values of $q/j^{2}$, the maximal frequencies of the modes increase with $j$ within the considered range. For large values of $q/j^{2}$, these frequencies increase only minimally or even decrease with increasing $j$.
Figure 5: Correlations between the lower and upper QPOs’ frequencies, $\nu_{\mathrm{L}}$ and $\nu_{\mathrm{U}}$, predicted by the CT model vs. data of the 4U 1636-53 atoll source. Non-rotating and moderately rotating NSs are assumed. Examples of the best fits predicted by the RP model for a given $j$ and $q$ are shown as well.
8.2 Consequences for models of NS variability
@@
Török et al. (2016a) and Török et al. (2022) have introduced a QPO model that identifies the twin peak QPO frequencies with the Keplerian frequency and the radial precession frequency of the fluid in the cusp-torus configuration (CT model). Within Figure 5, we compare the frequencies predicted by the model for two combinations of $M,j,q$, representing non-rotating and moderately rotating NSs, with the frequencies observed in the 4U 1636-53 atoll source (the data are taken from Barret et al., 2006; Török, 2009). We also include in the figure fits of the data based on the relativistic precession model (RP model, Stella & Vietri, 1999).
@@
From Figure 5, we can see that the RP model, which deals with geodesic precession, provides less promising fits of the data than the CT model, while the effects associated with the NS rotation do not imply a significant improvement (see Török et al., 2012, 2016b, 2016a, 2022). One can also see that even when we restrict ourselves to values of the Hartle-Thorne spacetime parameters that are consistent with up-to-date models of neutron stars, no conceivable smooth monotonic curve can reproduce the data in a significantly better way compared to the CT model. In principle, the relations derived here, including those determining the vertical precession frequency, can be used in an analogous way to put constrains on models of low-frequency QPOs.
@@
Figure 6: Left: Profiles of the slender torus vertical precession frequency for a particular choice of spacetime parameter’s values, along with the related NS radii calculated using L NS EoS (Pandharipande & Smith, 1975) and denoted by vertical lines. These are denoted by the vertical dashed lines. The top (bottom) panel corresponds to a rotational frequency of 400Hz (500Hz). Right: The range of the Hartle-Thorne spacetime parameters for which the vertical precession may change its orientation. The thick black line marked as $\beta=0$ indicates the limiting NS parameters relevant for the case of a slender torus, which are the same as for the geodesic (thin disk) case explored by Tsang & Pappas (2016). The shaded region below the $\beta=0$ line determines the area where this effect may occur, but the corresponding tori overlaps with the star’s surface. The coloured region above the $\beta=0$ line corresponds to the tori that do not overlap with the star, while the colour scale indicates the radii of their centres for the slender torus case. In a full analogy, the lines for $\beta=0.1$ and $\beta=0.2$ show the limiting parameters where the torus of the corresponding thickness would touch the star’s surface. In these cases, the corresponding $r_{0}$ colourmaps would be slightly different but almost identical to the $\beta=0$ map shown in this figure.
@@ 8.3 Impact of NS EoS
An example of the impact of the NS EoS choice can be demonstrated by the effect of the vertical precession’s change of orientation discussed in Section 6.3. In the left panel of Figure 6, we illustrate profiles of the vertical precession frequency for a particular choice of spacetime parameter’s values along with several related NS radii which are calculated using L NS EoS (Pandharipande & Smith, 1975). From the figure, we can see that the investigated effect occurs at the radii that are not necessarily above the NS surface.
In the right panel of Figure 6, we show the full range of NS’s angular momentum and quadrupole moment, which corresponds to the presence of the effect. Utilizing the recently identified universal relations (Maselli et al., 2013; Urbanec et al., 2013; Yagi & Yunes, 2013a, b; Pappas, 2015; Reina et al., 2017), we denote the coloured area which determines the range limited by the condition that the inner edge of the torus should be located above the NS surface. Clearly, this range is smaller than the full range given by the general consideration of the Hartle-Thorne metric alone.

### v3

@@ 6.1 Axisymmetric modes
The left panel of Fig. 1 provides a comparison between the axisymmetric ($m=0$) radial mode frequencies for tori with different thickness and NSs with $j=0.2$ and different compactness. The frequencies behave qualitatively the same as the test particle’s radial epicyclic frequency, being zero at ISCO, rising rather quickly with increasing radius to a distinct maximum and then slowly, almost steadily, decreasing. Because the pressure correction $\nu_{\mathrm{R}}^{(2)}$ is negative, the pressure forces tend to slow the epicyclic motion, and $\beta_{\mathrm{cusp}}$ rises with $r_{0}$, the maximum is located closer to the star than it would be in the case of a test particle epicyclic motion. A more detailed quantitative comparison is provided in Appendix A, Fig. 7.
In a full analogy to the left panel of Fig. 1, the right panel provides a comparison between the axisymmetric ($m=0$) vertical mode frequencies for tori with different thickness and NSs with $j=0.2$ and different compactness. Again, the frequencies behave qualitatively like a test particle’s vertical epicyclic frequency while the pressure corrections are negative. A more detailed quantitative comparison is also provided in Appendix A, Fig. 8.
@@ 6.2 Radial precession ($m=-1$ radial mode)
The left panel of Fig. 2 provides a comparison between the radial precession mode frequencies for tori with different thicknesses and NSs with $j=0.2$ and different compactness. For the radii close to ISCO, the pressure corrections are negative. The maximal precession frequency is achieved at $r_{0}=r_{\mathrm{ISCO}}$, where only infinitely slender torus with zero pressure corrections can exist. The maximum achievable frequency slightly drops with rising NS quadrupole moment as the ISCO moves to higher radii, but it is still higher than the precession rate for lower $q$ at the same (coordinate) radius.
For more distant radii, the corrections are smaller but can be positive, and at radii where they change the sign, the frequency does not depend on the torus thickness. This feature is present for both rotating and non-rotating stars. Radii relevant to this effect are shown in the right panel of Fig. 2. A more detailed quantitative comparison between the radial precession mode frequencies is provided in Appendix B, Fig. 9.
@@ 6.3 Vertical precession ($m=-1$ vertical mode)
The left panel of Fig. 3 provides a comparison between the vertical precession mode frequencies for tori with different thicknesses and NSs with $j=0.2$ and different compactness. The frequency behaviour is even more complex than in the case of radial precession.
For compact stars, the maximal frequencies are achieved at the ISCO. For the test-particle motion and large NS oblateness, the combined influence of the oblate shape of the star and frame-dragging induce non-monotonic frequency profiles with maximum above the ISCO and emergence of a zero-frequency radius where the precession changes its orientation (Kluźniak & Rosińska, 2013, 2014; Tsang & Pappas, 2016; Urbancová et al., 2019). The same happens for slender and non-slender tori. At the same time, the correction to the mode frequency exhibits a behaviour similar to the case of radial precession, and radii, where the correction changes its sign exists. The relevant radii are shown in the right panel of Fig. 3.
A more detailed quantitative comparison between the vertical precession mode frequencies is provided in Appendix B (Fig. 10), while an extended discussion on the mode frequency behaviour is given in Sect. 8.4.
@@ 7 Approximate relations
The formulae for the oscillation and precession frequencies investigated in the previous section are presented in our public codes. Their explicit forms are rather lengthy and potentially too complicated for practical calculations. Inspired by our previous work (Török et al., 2022), we provide their approximate versions that enable more efficient implementation.
For each of the four modes, we postulate an approximate formula that in the limit of a slender torus provides frequencies of free test particle motion, and in the limit of a marginally overflowing (cusp) torus, fits well the results of the exact calculations of the mode frequency. The numerical coefficients in the approximate formula are found by comparing them with the results of the exact mode frequency calculations, which are performed using the least squares method. Expressions that allow for calculations considering tori of general thickness are also included.
@@ 7.1 Axisymmetric radial mode
we obtain a relation for the axisymmetric radial oscillations of a torus with a cusp. The specific $\mathcal{R}$ function takes values between 0 and 1. Its dependence on spacetime parameters is illustrated in Fig. 4a while a comparison between the (exact) mode frequencies and the approximate relation is shown in Fig. 4b assuming very slowly rotating stars. A thorough comparison between the outcomes of exact calculations and the application of the approximate relation across the whole considered range of parameters is then provided in Appendix C (Fig. 11).
@@ 7.3 Radial precession (non-axisymmetric radial mode)
The radial precession frequency can be approximated as (Török et al., 2022)
@@
Choosing a constant $\mathcal{B}$, $\mathcal{B}=1$, our relation with a very high accuracy provides the value of the periastron precession frequency of test particle motion around oblate neutron stars.
@@
the resulting relation well matches the numerical findings for the cusp torus’s radial precession. The specific $\mathcal{B}$ function does not depend on the Keplerian frequency (radial coordinate $r$), unlike the case of the axisymmetric modes. For slowly rotating stars it takes values close to $\mathcal{B}=0.75$. The exact values are illustrated in Fig. 4e while a comparison between the (exact) mode frequencies and the approximate relation is shown in Fig. 4f assuming very slowly rotating stars. A thorough comparison between the outcomes of the exact calculations and the application of the approximate relation across the whole considered range of parameters is then provided in Appendix C (Fig. 13).
@@ 7.4 Vertical precession
$+\frac{9}{\nu_{0}}(q-j^{2})(\mathcal{V}_{0}^{-2/3}-7.2),$
we can describe the vertical precession of a cusp torus. The specific $\mathcal{T}$ function takes values very close to 1. The exact values are illustrated in Fig. 4g while a comparison between the (exact) mode frequencies and the approximate relation is shown in Fig. 4h assuming very slowly rotating stars. A thorough comparison between the outcomes of the exact calculations and the application of the approximate relation across the whole considered range of parameters is then provided in Appendix C (Fig. 14).
@@ 7.4.2 Tori of arbitrary thickness
To describe a torus of arbitrary thickness, we use
@@ 8 Discussion and conclusions
The impact of the torus thickness, in most cases, leads to relatively high negative corrections to the modes’ frequencies, which should be reflected within possible astrophysical applications. Furthermore, as found in Sects. 6.2 and 6.3, for the precession modes, the pressure corrections can be positive within a limited range of radii because the special radii exist where the precession frequency does not depend on the torus thickness. This can have observable consequences. Nevertheless, since we are constrained by the adopted perturbative approach and the effect manifests itself on a relatively large radius where the scatter between the frequencies calculated for different tori is generally very small, it will be useful to verify its presence by a suitable future numerical solution of Eq. (8).
Figure 5: Left: Correlations between the lower and upper QPO frequencies, $\nu_{\mathrm{L}}$ and $\nu_{\mathrm{U}}$, predicted by the CT and RP models, compared to the observed data from 4U 1636-53 atoll source. The best fits predicted for non-rotating NSs are denoted by thick solid lines, while the best fits for moderately rotating NSs with $j=0.22$ and $q=6j^{2}$ are denoted by dashed lines. The thin solid lines are then drawn for these values of $M$ and $j$ but different values of $q$ ($q=1.5j^{2}$). For the CT model, the light-coloured region emphasises the change of the frequencies predicted by the model with changing $q$. Middle: Relative change of the cusp torus precession frequencies with angular momentum at the radius where $\nu_{\mathrm{RP}}=0.75\nu_{\mathrm{K}}$. Within the RP and CT models, this radius corresponds to the bottom left group of data points shown in the left panel. The light-coloured region then emphasises the corresponding change of $\nu_{\mathrm{RP}}$. Right: Relative change of precession frequencies with the quadrupole parameter. Frequencies are calculated at the same radius as considered within the middle panel.
@@ 8.1 Influence of NS parameters
Both the NS rotation and rotationally induced oblateness affecting the angular momentum and quadrupole moment of the NS have significant effects on the frequencies of the investigated tori oscillations. The impact of the quadratic relations behind the nature of the modes is well illustrated in terms of quadrupole parameter $q/j^{2}$ used in our study, which for a NS with a given EoS and mass does not depend on the NS rotational frequency (Hartle & Thorne, 1968; Urbancová et al., 2019) . This parameter effectively ranging from $q/j^{2}\sim 1.5$ to $q/j^{2}\sim 10$ determines whether the influence of the $t$-$\phi$ component of the metric tensor and related frame dragging effects dominate (low values of the parameter corresponding to Kerr like behaviour) or are overwhelmed by the other components (high values of the parameter corresponding to the effective suppression of frame-dragging effects by the NS oblateness).
This interaction between the relativistic frame dragging effects associated with the angular momentum and the effects associated with the quadrupole moment, which would also arise in Newtonian physics, determines the resulting behaviour of the frequencies of the studied modes. For low values of $q/j^{2}$, the maximal frequencies of the modes increase with $j$ within the considered range. For large values of $q/j^{2}$, these frequencies increase only minimally or even decrease with increasing $j$.
8.2 Consequences for models of NS variability - HF QPOs
@@
Török et al. (2016) and Török et al. (2022) have introduced a QPO model that identifies the twin peak QPO frequencies with the Keplerian frequency, $\nu_{\mathrm{K}}=\Omega_{0}/2\pi$, and the radial precession frequency, $\nu_{\mathrm{RP}}$, of the fluid in the cusp-torus configuration (CT model). In the left panel of Fig. 5, we compare the best fits to the frequencies observed in the 4U 1636-53 atoll source (the data are taken from Barret et al., 2006; Török, 2009) obtained using the CT model to the best fits based on the relativistic precession model (RP model, Stella & Vietri, 1999; Stella et al., 1999).
@@
From Fig. 5, we can see that the RP model, which deals with geodesic precession, provides less promising fits of the data than the CT model. The effects associated with the NS rotation do not bring significant improvement in comparison to the non-rotating case and rather change the best-fitting mass value because the fitting-parameter space is almost degenerate (Török et al., 2016, 2022). Moreover, when we restrict ourselves to values of the Hartle-Thorne spacetime parameters that are consistent with up-to-date models of NSs, or impose even tighter constraints, such as $j<0.3$, $M<2.3M_{\sun}$, $q/j^{2}<10$, no conceivable smooth monotonic curve can reproduce the data in a significantly better way compared to the CT model.
8.3 Consequences for models of NS variability - LF QPOs
In principle, the relations derived here, including those determining the vertical precession frequency, can be used in an analogous way to put constraints on models of low-frequency QPOs providing that they are associated with the vertical precession. Following this research direction, initiated in the context of the RP model, can be very promising. In the middle and right panels of Fig. 5, we consider a particular orbital radius where $\nu_{\mathrm{RP}}=0.75\,\nu_{\mathrm{K}}$. Within the RP and CT models, this radius roughly corresponds to a large part of the documented QPO detections ($r=6.75r_{\mathrm{G}}$ when $j=0$). These panels clearly illustrate that the vertical precession frequency of the cusp torus is more sensitive to angular momentum and quadrupole moment than the radial one. This is caused by the strong vertical change of the gravitational field in the vicinity of an oblate NS, which already affects the test particle motion (Morsink & Stella, 1999; Kluźniak & Rosińska, 2013) and then applies to the fluid motion in tori of any thickness.
In order to explore detailed consequences of the behaviour of $\nu_{\mathrm{LT}}$ in relation to QPO models and diagnostics of NS parameters, further investigation is needed. As shown in the middle panel of Fig. 5, the dependence of $\nu_{\mathrm{LT}}$ on the NS spin can (for both test particle and fluid motion) exhibit a clear maximum, which can be of astrophysical importance.
@@ 8.3 Consequences for models of NS variability - LF QPOs
Figure 6: Left: Profiles of the slender torus vertical precession frequency for a particular choice of spacetime parameter’s values, along with the related NS radii calculated using the L NS EoS (Pandharipande & Smith, 1975) and denoted by vertical dashed lines. The top (bottom) panel corresponds to a rotational frequency of 400Hz (500Hz). Right: The range of the Hartle-Thorne spacetime parameters for which the vertical precession may change its orientation. The thick black line marked as $\beta=0$ indicates the limiting NS parameters relevant for the case of a slender torus, which are the same as for the geodesic (thin disc) case explored by Tsang & Pappas (2016). The shaded region below the $\beta=0$ line determines the area where this effect may occur, but the corresponding tori overlaps with the star’s surface. The coloured region above the $\beta=0$ line corresponds to the tori that do not overlap with the star, while the colour scale indicates the radii of their centres for the slender torus case. In a full analogy, the lines for $\beta=0.1$ and $\beta=0.2$ show the limiting parameters where the torus of the corresponding thickness would touch the star’s surface. In these cases, the corresponding $r_{0}$ colourmaps would be slightly different but almost identical to the $\beta=0$ map shown in this figure.
@@ 8.4 Impact of NS EoS
An example of the impact of the NS EoS choice can be demonstrated by the effect of the vertical precession’s change of orientation discussed in Sect. 6.3. In the left panel of Fig. 6, we illustrate profiles of the vertical precession frequency for a particular choice of spacetime parameter’s values along with several related NS radii which are calculated using the L NS EoS introduced by Pandharipande & Smith (1975). From the figure, we can see that the investigated effect occurs at the radii that are not necessarily above the NS surface. We use the L NS EoS, as it is rather extremely stiff and therefore well allows us to illustrate the effect of a large quadrupole moment, but the effect can be relevant for many currently considered EoS.
In the right panel of Fig. 6, we show the full range of NS’s angular momentum and quadrupole moment, which corresponds to the presence of the effect. Utilising the recently identified universal relations (Maselli et al., 2013; Urbanec et al., 2013; Yagi & Yunes, 2013a, b; Pappas, 2015; Reina et al., 2017), we denote the coloured area which determines the range limited by the condition that the inner edge of the torus should be located above the NS surface. Clearly, this range is smaller than the full range given by the general consideration of the Hartle-Thorne metric alone.